OpenStudy (anonymous):

Each trip, a bus drops off more fans at a concert. On the first trip, the bus dropped off 30. On the 14th trip, which was the last one, the bus dropped off 95 fans. How many fans did the bus deliver to the concert in total? (2 points) Question 4 options: 1) 875 2) 813 3) 455 4) 423

OpenStudy (anonymous):

can you explain to me? @kiira_x3

OpenStudy (anonymous):

@kiiraa_x3

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

assuming this is an arithmetic series, the sum is given by : \[\dfrac{n}{2}(a+l) = \dfrac{14}{2}(30+95) = 875\]

OpenStudy (anonymous):

yay thank you

ganeshie8 (ganeshie8):

np